# Zeroth-order log-concave sampling

### v1

@@ 2.1 Mixing analysis
Let $X_{i}=x$ be a current point, and let $f_{x}$ be the failure probability of a stage. For any event $A$,
[…]
=$\underset{(i)}{=}\sum_{j=1}^{\infty}\int\pi^{X|Y=y}(A)\,\gamma_{h}(y-x)\,\mathrm{d}y\times f_{x}^{j-1}(1-f_{x})$
$=\int\pi^{X|Y=y}(A)\,\gamma_{h}(y-x)\,\mathrm{d}y=\int\pi^{X|Y=y}(A)\,(\delta_{x}P_{h})(y)\,\mathrm{d}y$
$=(\delta_{x}P_{h}Q_{h})(A)\,,$
where $(i)$ follows from Lemma 2.6. Thus, $X_{i+1}$ is distributed as $\pi_{i}^{X}P_{h}Q_{h}$ when $X_{i}\sim\pi_{i}^{X}$. ∎
2.2 Query complexity per iteration
Recall from [KVZ24] a concentration phenomenon for $\pi^{Y}=\pi^{X}*\gamma_{h}$.
Lemma 2.8.
For a convex body $\mathcal{K}\subset\mathbb{R}^{d}$ containing a unit ball,
$\pi^{Y}(\mathbb{R}^{d}\backslash\mathcal{K}_{\delta})\leq\exp\bigl{(}-\frac{\delta^{2}}{2h}+\delta d\bigr{)}\,,$
where $\mathcal{K}_{\delta}=\{x\in\mathbb{R}^{d}:d(x,\mathcal{K})\leq\delta\}$.
Taking $h=(2cd^{2})^{-1}$ and $\delta=1/d$ for some $c>0$ ensures $\pi^{Y}(\mathbb{R}^{d}\backslash\mathcal{K}_{\delta})\leq e^{-c+1}$. Note that the success probability of the rejection sampling at $Y_{k}=y$ is $\ell(y):=\mathbb{P}_{\mathcal{N}(y,hI_{d})}(\cdot\in\mathcal{K})$, so the failure probability of a stage in the $k$-th iteration is
$\int\bigl{(}1-\ell(y)\bigr{)}^{\tau}\,\mathrm{d}\pi_{k}^{Y}(y)=\mathbb{E}_{\pi_{k}^{Y}}[(1-\ell)^{\tau}]=:f_{k}\,.$
Let $n_{k}$ denote the number of trials required to complete the $k$-th iteration. Then, the total expected number of queries over $N$ iterations is $\sum_{k=1}^{N}\mathbb{E}n_{k}$. By the law of total expectation,
$\mathbb{E}n_{k}=\sum_{i=1}^{\infty}\mathbb{E}[n_{k}\,|\,\text{success in }i\text{-th stage}]\,\mathbb{P}(\text{success in }i\text{-th stage})\leq\sum_{i=1}^{\infty}i\tau\times f_{k}^{i-1}(1-f_{k})=\frac{\tau}{1-f_{k}}\,.$
Therefore, bounding the per-iteration complexity reduces to bounding $f_{k}$ for each $k\in[N]$.
Let $p$ be the conjugate of $q$ (i.e., $p^{-1}+q^{-1}=1$ for $q\geq 2$). Then,
$\mathbb{E}_{\pi_{k}^{Y}}[(1-\ell)^{\tau}]\leq M_{q}\,\bigl{(}\mathbb{E}_{\pi^{Y}}[(1-\ell)^{\tau p}]\bigr{)}^{1/p}\,,$
and
$\int_{\mathbb{R}^{d}}(1-\ell)^{\tau p}\,\mathrm{d}\pi^{Y}=\int_{\mathbb{R}^{d}\backslash\mathcal{K}_{\delta}}\cdot+\int_{\mathcal{K}_{\delta}\cap[\ell\geq\tau^{-1}\log(3M_{q})]}\cdot+\int_{\mathcal{K}_{\delta}\cap[\ell<\tau^{-1}\log(3M_{q})]}\cdot=:\mathsf{A}+\mathsf{B}+\mathsf{C}\,.$
Using $\pi^{Y}=\ell/\operatorname{vol}\mathcal{K}$ when bounding $\mathsf{C}$ below,
$\mathsf{A}$ $\leq\pi^{Y}(\mathbb{R}^{d}\backslash\mathcal{K}_{\delta})\leq e^{-c+1}\,,$
$\mathsf{B}$ $\leq\int_{[\ell\geq\tau^{-1}\log(3M_{q})]}\exp(-\ell\tau p)\,\mathrm{d}\pi^{Y}\leq\bigl{(}\frac{1}{3M_{q}}\bigr{)}^{p},$
$\mathsf{C}$ $\leq\int_{\mathcal{K}_{\delta}\cap[\ell<\tau^{-1}\log(3M_{q})]}\frac{\ell(y)}{\operatorname{vol}\mathcal{K}}\,\mathrm{d}y\leq\frac{\log(3M_{q})}{\tau}\,\frac{\operatorname{vol}\mathcal{K}_{\delta}}{\operatorname{vol}\mathcal{K}}\leq\tau^{-1}e\log(3M_{q})\,,$
where we used $\operatorname{vol}\mathcal{K}_{\delta}\subset\operatorname{vol}((1+\delta)\,\mathcal{K})=(1+\delta)^{d}\,\operatorname{vol}\mathcal{K}\leq e\operatorname{vol}\mathcal{K}$ (due to $B_{1}(0)\subset\mathcal{K}$). Taking $c\asymp\log(3M_{q})^{p}$ and $\tau\asymp M_{q}^{p}\log(3M_{q})$, we can ensure that $\mathsf{A}+\mathsf{B}+\mathsf{C}\leq(\nicefrac{{1}}{{2M_{q}}})^{p}$ and thus $f_{k}\leq\frac{1}{2}$.
Putting these together, we bound the total query complexity during $N$ iterations as
$\sum_{k=1}^{N}\mathbb{E}n_{k}\leq\sum_{k=1}^{N}\frac{\tau}{1-f_{k}}\leq 2\tau N\,.$
Using this, we derive the complexity of $\mathsf{PS}_{\textup{unif}}$ with balanced guarantees.
@@ 2.2 Query complexity per iteration
For $M:=\mathcal{R}_{q}(\pi_{0}^{X}\mathbin{\|}\pi^{X})$, set $h^{-1}\asymp d^{2}M$ and $\tau\asymp Me^{M}$. To achieve $\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq\varepsilon$, by Lemma 2.5, it suffices to iterate at most
$N\lesssim q\,(1\vee h^{-1}\Lambda\log d)\log\frac{M}{\varepsilon}\lesssim q\Lambda d^{2}M\log d\log\frac{M}{\varepsilon}\quad\text{times}\,.$
Therefore, the expected number of total queries during $N$ iterations is
$2\tau N\lesssim q\Lambda d^{2}M^{2}e^{M}\log d\log\frac{M}{\varepsilon}=\widetilde{O}\bigl{(}q\Lambda d^{2}e^{\mathcal{R}_{q}(\pi_{0}^{X}\mathbin{\|}\pi^{X})}\log\frac{1}{\varepsilon}\bigr{)}\,,$
which completes the proof. ∎
@@ 3.1 Adjoint of the heat semigroups and the proximal sampler
We first recap the Markov semigroup theory for a gentle introduction to heat semigroup and its adjoint. Interested readers may want to refer to [VH14, BGL14, Che25] for further details on the semigroup theory. Meanwhile, we take detour to $\mathsf{PS}$, since it sets up a context in which the introduction of heat semigroups and its adjoint is natural.
@@
$(P_{t}f)(x):=\mathbb{E}[f(X_{t})|X_{0}=x]\quad\text{for every bounded measurable function }f:E\to\mathbb{R}\,.$
@@
As examined in [CCSW22] (with further details in [Che25]), the forward step (i.e., $y_{k+1}\sim\pi^{Y|X=x_{k}}=\mathcal{N}(x_{k},hI_{d})$) can be viewed as simulating Brownian motion for time $h$ started at $x_{k}$; when $\mathrm{d}X_{t}=\mathrm{d}B_{t}$ with $X_{0}=x_{k}$, we have $X_{h}\sim\operatorname{law}y_{k+1}$. The associated Markov semigroup of Brownian motion is often called the heat semigroup, which admits the formula of $P_{t}f=f*\gamma_{t}=:f_{t}$. Based on the classical heat equation $\partial_{t}P_{t}f=\frac{1}{2}\,\Delta P_{t}f$, rich theory for the heat semigroup facilitates one to understand the forward step in a better way (e.g., contraction under simultaneous heat flow).
@@
$\overline{Q}_{h}f(y):=\mathbb{E}[f(Z_{h}^{\leftarrow})|Z_{0}^{\leftarrow}=y]=\mathbb{E}_{\pi^{X|Y=y}}f\,.$
@@
This heat adjoint could be viewed as the backward step of $\mathsf{PS}$ with $\pi^{X}=\pi$ and $t=h$, since
$\overline{Q}_{t}f(y)=\int f\,\mathrm{d}\pi^{X|Y=y}=\frac{\int f(x)\,\pi^{X}(x)\,\gamma_{t}(y-x)\,\mathrm{d}x}{\int\pi^{X}(x)\,\gamma_{t}(y-x)\,\mathrm{d}x}=\frac{(f\pi)*\gamma_{t}}{\pi*\gamma_{t}}=\frac{P_{t}(f\pi)}{\pi_{t}}=(Q_{t}f)(y)\,.$
It follows from this formula that the adjoint operator satisfies mass preservation ($Q_{t}\mathds{1}=\mathds{1}$) and positivity preserving ($Q_{t}f\geq 0$ if $f\geq 0$). Moreover, it is a contraction operator; for $f\in L^{2}(\pi)\backslash\{0\}$, by Cauchy-Schwarz
@@
and integration of both sides yields $\lVert Q_{t}f\rVert_{L^{2}(\pi_{t})}^{2}\leq\lVert f\rVert_{L^{2}(\pi)}^{2}$.
@@ 3.2 Hypercontractivity of an adjoint of the heat semigroups
[…]
@@
Just as in the SDPI under ($\mathsf{PI}$), the SDPI under ($\mathsf{LSI}$) was established in [CCSW22, KO25]: for $q>1$ and two probability measures $\mu,\nu$ with $\mu\ll\nu$,
@@
Therefore, $\mathsf{PS}$ suffices to iterate $N=N_{0}+N_{1}\lesssim h^{-1}C_{\mathsf{LSI}}(\pi)\log\frac{q\mathcal{R}_{q}(\pi_{0}\mathbin{\|}\pi)}{(1\wedge(p-1))\,\varepsilon}$ times to attain the desired guarantee. ∎
As a simple corollary, this result leads to a mixing time of $\mathsf{PS}$ in the well-conditioned setting, which improves the linear dependence of $\mathsf{PS}$ on $q$ [CCSW22].
@@
4 Order-preserving annealing for simpler warm-start generation
In this section, we streamline prior warm-start generation algorithms and their proofs. Above all, we interweave an overall analysis with hypercontractivity under the heat flow, eventually establishing $\mathcal{R}_{q}$-warmness.
@@ 4 Order-preserving annealing for simpler warm-start generation
    Sample from $\mu_{1}=\pi\gamma_{d^{-1}}$ by rejection sampling with proposal $\gamma_{d^{-1}}$.
@@
    While $\sigma_{i}^{2}\lesssim qR\Lambda^{1/2}\log^{1/2}d$, move from $\bar{\mu}_{i}$ to $\bar{\mu}_{i+1}$ by $\mathsf{PS}_{\textup{Gauss}}$ such that $\mathcal{R}_{3}(\bar{\mu}_{i}\mathbin{\|}\mu_{i})=O(1)$, where
@@
We need two ingredients to obtain a final guarantee: (1) the query complexity of $\mathsf{PS}$ for truncated Gaussians and (2) closeness of two annealing distributions. The former was already studied in previous work, but it will be further simplified in §4.1 as done for uniform sampling in an earlier section. The latter was proven for $\chi^{2}$-divergence by a localization argument [CV18] and then was extended to $\mathcal{R}_{q}$ in [KV25a]. We show in §4.2 that this is in fact a simple consequence of ($\mathsf{LSI}$). Combining these together, we establish a query complexity of generating $\mathcal{R}_{q}$-warmness.
@@ 4.1 Proximal sampler for truncated Gaussian
We study $\mathsf{PS}$ for $\mu^{X}:=\pi\gamma_{\sigma^{2}}$, Gaussian distributions truncated to a convex body $\mathcal{K}$ containing $B_{1}(0)$. Its prototype was already studied in [KZ25, KV25a]. The streamlined algorithm, denoted by $\mathsf{PS}_{\textup{Gauss}}$, can be readily deduced from the general framework of $\mathsf{PS}$.
@@
[…]
As in the uniform-sampling case, its analysis is split into two parts: a mixing analysis and per-iteration complexity. The per-iteration analysis will be proceeded in a similar fashion, but the mixing analysis can be improved due to the hypercontractivity of $\mathsf{PS}$.
@@
Per-step query complexity.
Denoting by $\ell(y)=\mathbb{P}_{\mathcal{N}(sy,shI_{d})}(\cdot\in\mathcal{K})$ the success probability of the rejection sampling, we can readily check that
$\mu^{Y}(y)=(\mu^{X}*\gamma_{h})(y)=\frac{s^{d/2}\ell(y)\exp(-\frac{1}{2s\sigma^{2}}\,|sy|^{2})}{\int_{\mathcal{K}}\exp(-\frac{1}{2\sigma^{2}}\,|x|^{2})\,\mathrm{d}x}\,,$
and also recall a concentration phenomenon of $\mu^{Y}$ from [KZ25]: for $R=s^{-1}\mathcal{K}_{\delta}$, if $\delta\geq hd$,
$\mu^{Y}(\mathbb{R}^{d}\backslash R)\leq\exp\bigl{(}-\frac{\delta^{2}}{2h}+\delta d+hd^{2}\bigr{)}\,.$
Taking $h=(cd^{2})^{-1}$ and $\delta=d^{-1}$ for some $c\geq 1$, we have $\mu^{Y}(\mathbb{R}^{d}\backslash R)\leq e^{-c/2+2}$. Just as we used the bound of $\operatorname{vol}\mathcal{K}_{\delta}/\operatorname{vol}\mathcal{K}\leq e^{\delta d}$, we need the following “expansion” lemma (see [KZ25, Lemma 24]):
$s^{-d/2}\int_{\mathcal{K}_{\epsilon}}\exp\bigl{(}-\frac{1}{2s\sigma^{2}}\,|z|^{2}\bigr{)}\,\mathrm{d}z\leq 2e^{hd^{2}+\epsilon d}\int_{\mathcal{K}}\exp\bigl{(}-\frac{1}{2\sigma^{2}}\,|z|^{2}\bigr{)}\,\mathrm{d}z\,.$
This leads to the following bound: by change of variables via $y=s^{-1}v$ below,
$\frac{\int_{R}s^{d/2}\exp(-\frac{1}{2s\sigma^{2}}\,|sy|^{2})\,\mathrm{d}y}{\int_{\mathcal{K}}\exp(-\frac{1}{2\sigma^{2}}\,|x|^{2})\,\mathrm{d}x}=\frac{\int_{\mathcal{K}_{\delta}}s^{-d/2}\exp(-\frac{1}{2s\sigma^{2}}\,|v|^{2})\,\mathrm{d}v}{\int_{\mathcal{K}}\exp(-\frac{1}{2\sigma^{2}}\,|v|^{2})\,\mathrm{d}v}\leq 2e^{hd^{2}+\delta d}\leq e^{3}\,.$ (4.1)
We show that the failure probability of one stage in the $k$-th iteration is at most $1/2$ under suitable choices of $h$ and $\tau$. For a conjugate pair $(p,q)$ with $p,q>1$, we have $\mathbb{E}_{\mu_{k}^{Y}}[(1-\ell)^{\tau}]\leq M_{q}\,\{\mathbb{E}_{\mu^{Y}}[(1-\ell)^{\tau p}]\}^{1/p}$ and
$\int_{\mathbb{R}^{d}}(1-\ell)^{\tau p}\,\mathrm{d}\mu^{Y}=\int_{\mathbb{R}^{d}\backslash R}\cdot+\int_{R\cap[\ell\tau\geq\log(3M_{q})]}\cdot+\int_{R\cap[\ell\tau<\log(3M_{q})]}\cdot=:\mathsf{A}+\mathsf{B}+\mathsf{C}\,.$
Under $c\asymp\log(3M_{q})^{p}$ and $\tau\asymp M_{q}^{p}\log(3M_{q})$, this can be bounded by $(\nicefrac{{1}}{{2M_{q}}})^{p}$ due to
$\mathsf{A}$ $\leq\mu^{Y}(\mathbb{R}^{d}\backslash\mathcal{K}_{\delta})\leq e^{-c/2+2}\,,$
$\mathsf{B}$ $\leq\int_{R\cap[\ell\tau\geq\log(3M_{q})]}e^{-\ell\tau p}\,\mathrm{d}\mu^{Y}\leq\bigl{(}\frac{1}{3M_{q}}\bigr{)}^{p}\,,$
$\mathsf{C}$ $\leq\int_{R\cap[\ell\tau<\log(3M_{q})]}\mathrm{d}\mu^{Y}=\int_{R\cap[\ell\tau<\log(3M_{q})]}\frac{s^{d/2}\ell(y)\exp(-\frac{1}{2s\sigma^{2}}\,|sy|^{2})}{\int_{\mathcal{K}}\exp(-\frac{1}{2\sigma^{2}}\,|x|^{2})\,\mathrm{d}x}$
[…]

### v2

@@ 2.2.2 Per-step analysis
The uncapped $\mathsf{PS}_{\textup{unif}}$ can ensure $\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq\varepsilon$ after iterating
$N\gtrsim q\,\bigl(1\vee h^{-1}C_{\mathsf{PI}}(\pi)\bigr)\log\frac{M_{q}}{\varepsilon}\,.$
Under the choice of $h=cd^{-2}=\frac{\log\log Z}{2d^{2}\log Z}$ with $Z\asymp\frac{NM_{2}}{\eta}$, the required number $N$ of iterations must satisfy an inequality of the form
$N\gtrsim A\log(BN)\,$,
which can be fulfilled if $N\gtrsim A\log(AB)$. Therefore, the required number of iterations is of order
$N=\widetilde{O}\bigl(qd^{2}C_{\mathsf{PI}}(\pi)\log\frac{M_{q}}{\varepsilon}\log\frac{M_{q}}{\eta}\bigr)=\widetilde{O}\bigl(qd^{2}\Lambda\log\frac{1}{\varepsilon}\log\frac{1}{\eta}\bigr)\,.$
where $\Lambda=\lVert\operatorname{cov}\pi\rVert$, and the last equality follows from the assumption $M_{q}\leq 10$.
We now find a sufficient condition for ensuring $q\geq 1+\log\tau$. Observe that when $q\geq 2$,
$1+\log\tau=1+\frac{q}{q-1}\log(Z\log^{4}Z)\lesssim\log Z\lesssim\log\frac{N}{\eta}\,.$
Hence, when $q\geq 2\vee O(\log\frac{N}{\eta})$, then expected total query complexity of $\mathsf{PS}_{\textup{unif}}:\mathcal{R}_{q}\to\mathcal{R}_{q}$ from an $O(1)$-warm start is
$\widetilde{O}\bigl(qd^{2}\Lambda\log\frac{1}{\varepsilon}\log\frac{1}{\eta}\times\log^{4}Z\bigr)=\widetilde{O}\bigl(qd^{2}\Lambda\log\frac{1}{\varepsilon}\log^{5}\frac{1}{\eta}\bigr)\,.$
Lastly, the success probability satisfies $\mathbb{P}(S)\geq\eta$ by construction, and by Lemma 2.7, the output law $\hat{\pi}_{N}^{X}$ conditioned on success satisfies
$\mathcal{R}_{q}(\hat{\pi}_{N}^{X}\mathbin{\|}\pi^{X})\leq\varepsilon+2\log\frac{1}{1-\eta}\,$,
which completes the proof. ∎
Using this proposition, we prove Theorem 1.1.
@@
Take $\eta=\varepsilon<0.01$, and restart the algorithm upon failure. Note that the expected number of restarts is at most $0.02$. To invoke the proposition above, $q$ should satisfy $q\geq 2\vee O(\log N)$ for $N=\widetilde{O}(qd^{2}\Lambda\log^{2}\frac{1}{\varepsilon})$. To ensure
$q\geq\log\widetilde{O}\bigl(qd^{2}\Lambda\log^{2}\frac{1}{\varepsilon}\bigr)(\geq\log N)\,$,
let us solve the inequality of the form $q\gtrsim A\log qB$. A sufficient condition for this is simply $q\gtrsim A\log AB=\widetilde{O}(A\log B)$, so it suffices to enforce
$q\geq 2\vee\widetilde{O}\bigl(\log d^{2}\Lambda+\log\log\frac{1}{\varepsilon}\bigr)\,.$ (2.3)
Hence, by Proposition 2.8, the restart version of $\mathsf{PS}_{\textup{unif}}$ uses $\widetilde{O}(qd^{2}\Lambda\log^{6}\frac{1}{\varepsilon})$ total queries in expectation, and its output is distributed as $\hat{\pi}_{N}^{X}$ satisfying $\mathcal{R}_{q}(\hat{\pi}_{N}^{X}\mathbin{\|}\pi^{X})\leq 3\varepsilon$. Retaking $\varepsilon\leftarrow\varepsilon/3$, we complete the proof. ∎
@@ 3.1 Adjoint of the heat semigroups and the proximal sampler
We first recap the Markov semigroup theory for a gentle introduction to heat semigroup and its adjoint. Interested readers may want to refer to [VAN14, BGL14, CHE25] for further details on the semigroup theory. Meanwhile, we take detour to $\mathsf{PS}$, since it sets up a context in which the introduction of heat semigroups and its adjoint is natural.
@@
$(P_{t}f)(x):=\mathbb{E}[f(X_{t})\mid X_{0}=x]\quad\text{for every bounded measurable function }f:E\to\mathbb{R}\,.$
@@
As examined in [CCS+22] (with further details in [CHE25]), the forward step (i.e., $y_{k+1}\sim\pi^{Y|X=x_{k}}=\mathcal{N}(x_{k},hI_{d})$) can be viewed as simulating Brownian motion for time $h$ started at $x_{k}$; when $\mathrm{d}X_{t}=\mathrm{d}B_{t}$ with $X_{0}=x_{k}$, we have $X_{h}\sim\operatorname{law}y_{k+1}$. The associated Markov semigroup of Brownian motion is often called the heat semigroup, which admits the formula of $P_{t}f=f*\gamma_{t}=:f_{t}$. Based on the classical heat equation $\partial_{t}P_{t}f=\frac{1}{2}\,\Delta P_{t}f$, rich theory for the heat semigroup facilitates one to understand the forward step in a better way (e.g., contraction under simultaneous heat flow).
@@
$\overline{Q}_{h}f(y):=\mathbb{E}[f(Z_{h}^{\leftarrow})\mid Z_{0}^{\leftarrow}=y]=\mathbb{E}_{\pi^{X|Y=y}}f\,.$
@@
This formula can be naturally derived, though not mentioned in the original manuscript, from the definition of adjoint: for any $f\in L^{2}(\pi)$ and $g\in L^{2}(\pi_{t})$,
$\langle f,P_{t}g\rangle_{L^{2}(\pi)}=\langle Q_{t}f,g\rangle_{L^{2}(\pi_{t})}\,.$
For $\mathrm{d}Z_{t}=\mathrm{d}B_{t}$, one could rewrite the LHS as follows:
$\langle f,P_{t}g\rangle_{L^{2}(\pi)}=\int f\,P_{t}g\,\mathrm{d}\pi=\mathbb{E}\bigl[f(Z_{0})\,\mathbb{E}[g(Z_{t})|Z_{0}]\bigr]=\mathbb{E}[f(Z_{0})\,g(Z_{t})]=\mathbb{E}_{Z_{t}\sim\pi_{t}}\bigl[g(Z_{t})\,\mathbb{E}[f(Z_{0})|Z_{t}]\bigr]\,.$
Equating this with $\langle Q_{t}f,g\rangle_{L^{2}(\pi_{t})}$, we can deduce that
$Q_{t}f(y)=\mathbb{E}[f(Z_{0})|Z_{t}=y]\,.$
Since $\operatorname{law}(Z_{0}|Z_{t}=y)=\pi^{X|Y=y}$ with $\pi^{X}=\pi$,
$Q_{t}f(y)=\int f\,\mathrm{d}\pi^{X|Y=y}=\frac{\int f(x)\,\pi^{X}(x)\,\gamma_{t}(y-x)\,\mathrm{d}x}{\int\pi^{X}(x)\,\gamma_{t}(y-x)\,\mathrm{d}x}=\frac{(f\pi)*\gamma_{t}}{\pi*\gamma_{t}}=\frac{P_{t}(f\pi)}{\pi_{t}}\,.$
We note that this heat adjoint could be viewed as the backward step of $\mathsf{PS}$: $\overline{Q}_{t}f=Q_{t}f$. Also, it follows from the above formula that the adjoint operator satisfies mass preservation ($Q_{t}\mathds{1}=\mathds{1}$) and positivity preserving ($Q_{t}f\geq 0$ if $f\geq 0$). Moreover, it is a contraction operator; for $f\in L^{2}(\pi)\backslash\{0\}$, by Cauchy-Schwarz
@@
and integration of both sides yields $\lVert Q_{t}f\rVert_{L^{2}(\pi_{t})}^{2}\leq\lVert f\rVert_{L^{2}(\pi)}^{2}$. We refer readers to [KV25c, §B.2].
@@ 3.2 Hypercontractivity of an adjoint of the heat semigroups
$\bar{f}_{t}^{q^{\prime}}\log\bar{f}_{t}=(\bar{f}_{t}\,\mathds{1}[\bar{f}_{t}\geq 1]+\bar{f}_{t}\,\mathds{1}[\bar{f}_{t}<1])^{q^{\prime}}\log\bar{f}_{t}\leq 2^{q+\delta-1}\,(\bar{f}_{t}^{q+\delta}+\bar{f}_{t}^{q-\delta})\log\bar{f}_{t}\,.$
@@
Just as in the SDPI under ($\mathsf{PI}$), the SDPI under ($\mathsf{LSI}$) was established in [CCS+22, KO25]: for $q>1$ and two probability measures $\mu,\nu$ with $\mu\ll\nu$,
@@
Therefore, $\mathsf{PS}$ suffices to iterate $N=N_{0}+N_{1}\lesssim h^{-1}C_{\mathsf{LSI}}(\pi)\log\frac{q\mathcal{R}_{2\wedge p}(\pi_{0}\mathbin{\|}\pi)}{(1\wedge(p-1))\,\varepsilon}$ times to attain the desired guarantee. ∎
As a simple corollary, this result leads to a mixing time of $\mathsf{PS}$ in the well-conditioned setting, which improves the linear dependence of $\mathsf{PS}$ on $q$ [CCS+22].
@@
4 Order-preserving annealing via simpler warm-start generation
In this section, we streamline prior warm-start generation algorithms and their proofs. Above all, we interweave an overall analysis with hypercontractivity under the heat flow, eventually establishing the complexity of generating $\mathcal{R}_{q}$-warmness.
@@ 4 Order-preserving annealing via simpler warm-start generation
    Sample from $\mu_{1}=\pi\gamma_{d^{-1}}$ by rejection sampling with proposal $\gamma_{d^{-1}}$, and set $\bar{\mu}_{1}:=\mu_{1}$.
@@
    While $\sigma_{i}^{2}\lesssim qR\Lambda^{1/2}\log^{1/2}d$, move from $\bar{\mu}_{i}$ to $\bar{\mu}_{i+1}$ by $\mathsf{PS}_{\textup{Gauss}}$ such that $\mathcal{R}_{\widetilde{O}(1)}(\bar{\mu}_{i+1}\mathbin{\|}\mu_{i+1})=O(1)$, where
@@
We need two ingredients to obtain a final guarantee: (1) the query complexity of $\mathsf{PS}$ for truncated Gaussians ($\mathsf{PS}_{\textup{Gauss}}$) and (2) closeness of two annealing distributions. The former was already studied in previous work, but we will refine the analysis to obtain an order-boosting guarantee via hypercontractivity in §4.1. The latter was proven for $\chi^{2}$-divergence by a localization argument [CV18] and then was extended to $\mathcal{R}_{q}$ in [KV25a]. We show in §4.2 that this is in fact a simple consequence of ($\mathsf{LSI}$). Combining these together, we establish the query complexity of generating $\mathcal{R}_{q}$-warmness.
@@ 4.1 Proximal sampler for truncated Gaussian
We study $\mathsf{PS}$ for $\mu^{X}:=\pi\gamma_{\sigma^{2}}$, Gaussian distributions truncated to a convex body $\mathcal{K}$ containing $B_{1}(0)$. Its prototype was already studied in [KZ25, KV25a], and $\mathsf{PS}_{\textup{Gauss}}$ can be readily deduced from the general framework of $\mathsf{PS}$.
@@
  • [Backward] $x_{k+1}\sim\mu^{X|Y=y_{k+1}}\propto\gamma_{sh}(\cdot-sy_{k+1})\,\mathds{1}_{\mathcal{K}}(\cdot)$. To this end, draw $x_{k+1}\sim\mathcal{N}(sy_{k+1},shI_{d})$ repeatedly until $x_{k+1}\in\mathcal{K}$. If this rejection loop exceeds $\tau$ trials, declare failure and restart from scratch with a new sample $x_{0}\sim\mu_{0}^{X}$.
As in the uniform-sampling case, its analysis is split into two parts: a mixing analysis and per-iteration complexity. The failure probability and per-iteration analysis will be proceeded in a similar fashion, but the mixing analysis can be improved due to the hypercontractivity of $\mathsf{PS}$.
@@
Per-step analysis.
The analysis is analogous to §2.2, with $\pi_{k}^{Y}$ replaced by $\mu_{k}^{Y}$ and §2.1 of [KV25a] replaced by §2.2. With the parameter choices $Z=\frac{16NM_{2}}{\eta}$, $c=\frac{\log\log Z}{10\log Z}$, $h=\frac{c}{d^{2}}$, $\tau=Z^{2}\log^{3}Z$, the failure probability satisfies $\mathbb{E}_{\mu_{k}^{Y}}[(1-\ell)^{\tau}]\leq\frac{\eta}{N}$, and enforcing $q\geq 1+(1\vee\log\tau)$ yields $\mathbb{E}_{\mu_{k}^{Y}}[\frac{1}{\ell}\wedge\tau]\leq 5M_{q_{0}}\log^{4}Z$.
Proposition 4.1.
Let $\mathcal{K}\subset\mathbb{R}^{d}$ be a convex body with $B_{1}(0)\subset\mathcal{K}$, given access to a membership oracle for $\mathcal{K}$, and $\mu\propto\gamma_{\sigma^{2}}\cdot\mathds{1}_{\mathcal{K}}$ be the Gaussian truncated to $\mathcal{K}$. Given $\varepsilon>0$ , $\eta\in(0,1/2)$, $q\geq q_{0}\geq 2$, and an initial distribution $\mu_{0}$ with $M_{q_{0}}=\lVert\nicefrac{{\mathrm{d}\mu_{0}}}{{\mathrm{d}\mu}}\rVert_{L^{q_{0}}(\mu)}$, initialize $\mathsf{PS}_{\textup{Gauss}}$ with $\mu_{0}$ and iterate it $N=\widetilde{O}(d^{2}\sigma^{2}\log\frac{q}{\varepsilon}\log\frac{1}{\eta})$ times with $h=(10d^{2}\log Z)^{-1}$ and $\tau=Z^{2}\log^{3}Z$ for $Z=\frac{16NM_{2}}{\eta}$. If $q_{0}\geq 1+\log\tau$ and $M_{q_{0}}\leq 10$, then
  • The sufficient condition on $q_{0}$ is satisfied when $q_{0}\geq 2\vee O(\log\frac{N}{\eta})$.
  • With probability at least $1-\eta$, $\mathsf{PS}_{\textup{Gauss}}$ iterates $N$ times without failure. Conditioned on this success, the output $Z\sim\nu$ satisfies $\mathcal{R}_{q}(\nu\mathbin{\|}\mu)\leq\varepsilon+2\log\frac{1}{1-\eta}$, with $\widetilde{O}(d^{2}\sigma^{2}\log\frac{q}{\varepsilon}\log^{5}\frac{1}{\eta})$ membership queries in expectation.
Proof.
For $q\geq q_{0}\geq 2$ and $(M_{2}\leq)M_{q_{0}}\leq 10$, the uncapped $\mathsf{PS}_{\textup{Gauss}}$ (with the choice of $h=cd^{-2}=\frac{\log\log Z}{10d^{2}\log Z}$ and $Z\asymp NM_{2}/\eta$) returns a sample $X_{N}$ such that $\mathcal{R}_{q}(\operatorname{law}X_{N}\mathbin{\|}\mu^{X})\leq\varepsilon$ after iterating
$h^{-1}\sigma^{2}\log\frac{q\log M_{2}}{\varepsilon}\quad\text{times}\,.$
Repeating the similar argument as in the uniform sampling case, the required number of iterations is of order
$N=\widetilde{O}\bigl(d^{2}\sigma^{2}\log\frac{q}{\varepsilon}\log\frac{1}{\eta}\bigr)\,.$ (4.1)
A sufficient condition for ensuring $q_{0}\geq 2\vee\log\tau$ is simply
$1+\log\tau\lesssim\log\frac{N}{\eta}\lesssim\log\widetilde{O}\bigl(\frac{d^{2}\sigma^{2}}{\eta}\log\frac{q}{\varepsilon}\log\frac{1}{\eta}\bigr)\,$, (4.2)
Under this condition, the per-step complexity is $O(\log^{4}Z)$, so the expected number of membership queries for $\mathsf{PS}_{\textup{Gauss}}:\mathcal{R}_{q_{0}}\to\mathcal{R}_{q}$ in one attempt is
$\widetilde{O}\bigl(d^{2}\sigma^{2}\log\frac{q}{\varepsilon}\log^{5}\frac{1}{\eta}\bigr)\,$,
and the success probability satisfies $\mathbb{P}(S)\geq\eta$. As in the analysis of $\mathsf{PS}_{\textup{unif}}$, the conditional law of the final output is distributed as $\hat{\mu}_{N}^{X}$ satisfying $\mathcal{R}_{q}(\nu\mathbin{\|}\mu)\leq\varepsilon+2\log\frac{1}{1-\eta}$. ∎